Hello
with this problem i'm suppose to use the slope intercept method to graph the equation 5 = 2/3x - y i'm to write the slope (m) and the intercept (b) on graph paper to draw a graph of this problem. Can somebody please help me write a solution to this problem? Thanks!!

Answer by wuwei96815(245)

First you have to put the equation in the slope-intercept form which is y=mx+b, where m = the slope and b= the y-intercept.

5=2/3x-y
5+y=2/3x
y = 2/3x-5

Therefore, the slope of the line generated by this equation will be 2/3 and the y-intercept will be -5.

You can find the x-interept for this equation by solving it for x and then finding the value of x when y=0.